OpenStudy (anonymous):

I start with 9-2x widgets. Each day, I produce 2x+1 widgets, and destroy 3-x widgets. After three days, how many widgets am I left with?

OpenStudy (mokeira):

after 3 days you produce 3(2x+1)=6x+3 she destroys 3(3-x)=9-3x (9-2x)+(6x+3)-(9-3x) 9-2x+6x+3-9+9x 9+3-9-2x+6x+9x 3+13x hope you have understood

OpenStudy (anonymous):

it says it's incorrect

OpenStudy (mokeira):

then it must be...lol. let me try again

OpenStudy (anonymous):

ok thanks

OpenStudy (mokeira):

i wen wrong somewhere i replaced 3 with 9 (9-2x)+(6x+3)-(9-3x) 9-2x+6x+3-9+3x 9+3-9-2x+6x+3x 3+7x

OpenStudy (mokeira):

is that right?

OpenStudy (anonymous):

oh thanks so much your so awesome !

OpenStudy (anonymous):

it's so right

OpenStudy (mokeira):

yaaaay! glad to help :)